This year, Sigma Inc. generated $636,000 income from its routine business operations. In addition, the corporation sold the following assets, all of which were held for more than 12 months: Initial Basis Acc. Depr.* Sale Price Marketable securities $ 157,600 $ 0 $ 74,000 Production equipment 115,600 92,480 36,000 Business realty: Land 235,500 0 241,750 Building 263,000 78,900 218,000 *Through date of sale. Required: Compute Sigma’s taxable income assuming that it used the straight-line method to calculate depreciation on the building and has no nonrecaptured Section 1231 losses. Recompute taxable income assuming that Sigma sold the securities for $174,000 rather than $74,000.

Note: I have part B correct with $705,430. It's part A I am having trouble with
Answer #1

Answer:- 668605

a)Ordinary gain :-

Production equipment = 36000 - (115600-92480)

= 12880

Total = 12880 + (25%×78900) = 32605

Ordinary gain :-

Buisness building = 218000-(263000 -78900)

= 33900

= 33900 - (78900×25%) = 14175

Business land = 241750-235500=6250

Total = 20425

Taxable income = 636000 + 32605 +20425 - 20425 *(74000 - 157600 ) take lower

= 668605

b)

636000 + 32605 + 20425 +174000 - 157600

= 705430